OpenStudy (anonymous):

Consider the following Qoutient: 12 11 * 10 * 9 * 8 * 7 *6 * 5 * 4 * 3 * 2 / 2^x (its 12 times 11 times 10 and so on... and "x" is in mid air after 2) What is the maximum value of x such that 2^x is divisible into the product shown? (Hint: How many two's can divide evenly into the product shown in the numerator? Start "cancelling" two's)

OpenStudy (anonymous):

...its not that hard guys..

OpenStudy (anonymous):

12 = 2*2*3 (running count = 2) 11 = prime 10 = 2*5 (3) 9 = not divisible by 2 8 = 2*2*2 (6) 7 = prime 6 = 2*3 (7) 5 = prime 4 = 2*2 (9) 3 = prime 2 = 2 (10) x = 10?